differentiate f(x)=e^((1/2)x+cosx)
1
Expert's answer
2011-10-07T08:28:02-0400
(e^((1/2)x+cosx))'=(e^((1/2)x+cosx))* (1/2)x+cosx)'=(e^((1/2)x+cosx))*(1/2-sinx)